Projection Admin 9/22



To do this evening/ was completed: See T6 issues in maintenance. 

Headset device out in T5 - William

Presentation Issues:

Preshow Clips/Playlist: Tech'd Anomalisa (brought volume up and added a house lights down cue) and checked out Stolen Kingdom, which are both ready to go. - Lillian 

Trailer Updates:

Special Events/ AV set up: Checked in with Chenelle about content for the private event on 9/29, which she said should be locked down by tomorrow, Wednesday at the latest. - Lillian

Cristina let me know that the Gotham Slide should be coming to us this week and we're still waiting on them to confirm what kinda preshow they actually want. - Lillian 

Ordering/Receiving (posters & trailers):

Ordering/Receiving (content): The last two NewFest shorts came in today, which I've currently got downloading to the share drive. We've got our first tech this week! Woohoo! - Lillian 

Outgoing/Shipping (content): They Live will get picked up Wednesday. Sisterhood needs to go out, but we need a replacement reel for the one that came to us already broken. I've reached out about this. - Lillian 

Print/video Inspection:

Asset tracking doc updated/checked?:

Maintenance/cleaning: Got walkied about the TVs in Trees not working. After troubleshooting, checking connections, swapping out cables, etc, I have come to the conclusion that the av distribution component (splitter) is dying out (second option is that the hdmi to av is busted). Apple TV would briefly flash when I'd push the power cable in and out. Alec is aware of all of this and he's gonna check out what component Williamsburg uses since theirs has lasted ages. - Lillian 

More on the maintenance side - this morning we found that T6 was not "unreachable" on the LMS. I tried both powercycling the projector, IMS, and reseating the ethernet cable whilst the project was off and on, to no avail. Did a reboot of the LMS (which you can only do on the rack) and this did the trick. Logs and the drm report sent to DMS, which Gregg and their team have seen. Right now they'd like to update the firmware, but we'll hear more soon, hopefully. We'll keep you updated. Shows are running normally, but keep an eye on it. - Lillian 

Swept/mopped/cleaned booths 1 & 2 & ran loops for the show tonight - Will

Cue sheets updated/checked (every shift): Yes. 

Marquee updated?: Yes

Posters updated?:

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Open 12/28

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will